OpenStudy (anonymous):

The value of y varies directly with x2, and y = 64 when x = 4. What is the value of y when x = 6?

OpenStudy (anonymous):

if that x2 stands for \(x^2\), then: \[y=Ax^2\] where \(A\) is some constant, if you plug \(y=64\) for \(x=4\) you can find the value of \(A\) and then plug \(x=6\) and see what you get for \(y\)...

OpenStudy (anonymous):

6^2 would make it 36 right?

OpenStudy (anonymous):

@Greg_D

OpenStudy (anonymous):

that woul be the case \(A=1\)... you can et the value for \(A\) from: \[64=A\times 4^2\] find that and then calculate \(y=A\times 6^2\)

OpenStudy (anonymous):

what is a

OpenStudy (anonymous):

the problem says: "The value of y varies directly with x^2, and y = 64 when x = 4" so y must equal "something" times x^2, i called that something A

OpenStudy (anonymous):

im sorry im the worst XD

OpenStudy (anonymous):

576

OpenStudy (anonymous):

no problem, just keep asking until you understand ! what did you get for A ?

OpenStudy (anonymous):

well 64/4 is 16 so i multiplied 36 by 16 to get 576

OpenStudy (anonymous):

note that \(A=\frac{64}{4^2}\) dont forget the square!

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

i dont remember how to do that

OpenStudy (anonymous):

well \(4^2=16\) so \(A=\frac{64}{16}=4\)

OpenStudy (anonymous):

so 4 times 36 ?

OpenStudy (anonymous):

for 144?

OpenStudy (anonymous):

:) that is what i get... good work!